Order DSN 1,000 mLV at 50 ml/h for 20 hours. After 5 hours 700 mL remain to be infused. The infusion is ahead of schedule. Compute the new flow rate in mL/h so that the infusion will finish on time. Note that now rates may not be adjusted by more than 25% oftn original rate. To what rate would you reset the flow rate, or would you contact the prescriber? O A. Contact the prescriber O B. 40 mlh О С. 47 muh O D. 49 mh
media%2F65b%2F65b3ed7d-f8cf-4ae0-af63-bf
Answer #1

1. An IV is infusing at 750 mL/h. The strength of the solution in the bag is 10mcg/mL.

Thus, in 1 hour, 750*10 mcg = 7500 mcg = 7500*0.001mg = 7.5 mg would be infused ( as 1 mcg= 0.001mg).

Hence, the dosage rate is 7.5 mg/h.

2. Since 1000 mL is to be infused in 20 hours at 50 mL/h, hence in 5 hours, 5*50 = 250 mL should have been infused. However, the drug infusion is ahead of schedule and in 5 hours, 1000-700 = 300 mL of drug has been infused.

Thus, in 20-5 = 15 hours, 700 mL of the drug has to be infused . Hence the flow rate has to be adjusted at 700/15 = 46.67 or, 47 mL/h ( on rounding off to the nearest whole number) so that the infusion may finish on time. Option C is the correct answer.